OpenStudy (help_needed_cudi):

i need someone to check my questions and help me out with some if thats ok!! thanks!! i need help bfor 10! so please help

OpenStudy (help_needed_cudi):

is A 5squared 3

OpenStudy (help_needed_cudi):

oh in the last one is 15 but ik it wasnt that

OpenStudy (help_needed_cudi):

the next one is 4. Simplify the radical expression.

OpenStudy (help_needed_cudi):

i choose D

OpenStudy (help_needed_cudi):

5. Simplify the radical expression. (1 point) A. 245y²w³

OpenStudy (help_needed_cudi):

sorry this is answer C

OpenStudy (help_needed_cudi):

i dont know this question

OpenStudy (mathmale):

HNC: Please post one question at a time. It's difficult to know which of your five questions you're referring to. The key to solving the first problem is to look for a perfect square under the radical sign and then to separate (factor) the radical into two parts:\[\sqrt{45=}\sqrt{9}\sqrt{5}\] Is 9 a perfects quare? Is 5 a perfect square? What would you do next? What is your final answer?

OpenStudy (help_needed_cudi):

well no 9 is not so it would be 3 and 5 is.

OpenStudy (mathmale):

Let's back up a bit: Is 5 a perfect square?

OpenStudy (help_needed_cudi):

umm i think it is

OpenStudy (help_needed_cudi):

like it cant go any lower right?

OpenStudy (mathmale):

What is the square of 1?

OpenStudy (help_needed_cudi):

um nothing....so yes ig it can go all the way to 1

OpenStudy (mathmale):

In math, "squaring" a number involves multiplying it by itself. What is 1*1?

OpenStudy (help_needed_cudi):

1

OpenStudy (mathmale):

that's better. What are the squares of 2 and 3?

OpenStudy (help_needed_cudi):

1

OpenStudy (mathmale):

where did that come from? please go back to my explanation of the meaning of 'Squaring a number." The square of 2 is ... ??

OpenStudy (help_needed_cudi):

i had already answered that problem did i get it wrong or something? also i dont have alot of time and i really need help with the last ones that i havent answered.... 2

OpenStudy (help_needed_cudi):

ik but i wasnt sure so i didnt put it...

OpenStudy (mathmale):

I'm sorry, Cudi, but you have NOT answered Question 3 correctly, and you have not demonstrated that you understand what the "square" of a number is. Math is not a guessing game; you need to understand the definitions and the rules that apply. If you want to move on to the later questions, fine, but i will still expect you to invest enough time and effort so that you actually know what you 're doing. So...what do you want to do now?

OpenStudy (help_needed_cudi):

not guess btw....but thank you still for giving me your time and effort to help me i didnt answer number 3....

OpenStudy (mathmale):

Either you demonstrate how you obtain your answers or I must assume that you are guessing.

OpenStudy (mathmale):

Where is the "radical expression" of Problem #5?

OpenStudy (help_needed_cudi):

ok well for number 3. \[\sqrt{45} = \sqrt{5}*\sqrt{3} =5{\sqrt{3}}\]

OpenStudy (help_needed_cudi):

that is how it is in my math book so my math book must b wrong then...

OpenStudy (help_needed_cudi):

5.

OpenStudy (mathmale):

5 is NOT a perfect square, so Sqrt(5) does not equal 5. That's why I asked you to square the numbers 1, 2, 3, ... sorry, I do not agree that "my math book must be wrong." Plese decide which question you want to work on NOW.

OpenStudy (help_needed_cudi):

3....well 15 isnt either....im looking at the book and looking at yours and looking at the problems if its not then non of them are right...its not an argument its help so help me try to figure it out please and thank you.

OpenStudy (mathmale):

\[\sqrt{45} = \sqrt{9}*\sqrt{5} = ( ? )\sqrt{5}\]

OpenStudy (help_needed_cudi):

ohhh omg im embarrassed now i put the wrong thing in but yes i ment the 9 in the other one...3 squared of 5

OpenStudy (mathmale):

9 is a perfect square. What is the square root of 9? 5 is not a perfect square, so you cannot simplify Sqrt(5). Please type out your answer to #3.

OpenStudy (help_needed_cudi):

\[3\sqrt{5}\] the square root of 9 is 3

OpenStudy (mathmale):

\[\sqrt{45}=\sqrt{9}\sqrt{5}=3\sqrt{5}\] perfect!

OpenStudy (mathmale):

Which problem do you want to focus on next?

OpenStudy (help_needed_cudi):

number 5

OpenStudy (mathmale):

tYPE out #5 here, please, using Equation Editor.

OpenStudy (help_needed_cudi):

ok

OpenStudy (help_needed_cudi):

\[\sqrt{490}y ^{5} w ^{6} \]\[A. 245y ^{2} w ^{^{5}}\] \[B. 7y ^{2} w ^{3} \sqrt{10y}\]

OpenStudy (help_needed_cudi):

not done i have a very slow computer

OpenStudy (help_needed_cudi):

C.\[7 y ^{^{2}} w ^{3}\sqrt{10yw}\] D.\[7\sqrt{10y ^{5}} w ^{6}\]

OpenStudy (help_needed_cudi):

ok that took 4ever but there

OpenStudy (mathmale):

Is this the entire problem statement? \[\sqrt{490}y ^{5} w ^{6}\] If so, you don't have to do anything with the y^5 or the w^6. Instead, focus on Sqrt(490). There is a perfect square factor in 490. What is it?

OpenStudy (help_needed_cudi):

49?

OpenStudy (help_needed_cudi):

and yes that is the full statement

OpenStudy (mathmale):

What is the square root of 49?

OpenStudy (help_needed_cudi):

7

OpenStudy (mathmale):

All right. Then your entire expression simplifies to \[7\sqrt{10}y^5w^6.\]

OpenStudy (help_needed_cudi):

ok thanks again could you do one more?

OpenStudy (help_needed_cudi):

6. Simplify the radical expression.

OpenStudy (help_needed_cudi):

if possible?

OpenStudy (mathmale):

type out the last problem and do as much as YOU can do by YOURSELF. Then I'll respond as appropriate.

OpenStudy (help_needed_cudi):

ok

OpenStudy (mathmale):

All you need to share is the expression you need to simplify. Don't bother with the answers.